# Homomorphism counting for immersion-closed classes is not isomorphism

### v1

@@ Abstract
Lovász proved that two graphs $G$ and $H$ are isomorphic if $\hom(K,G)=\hom(K,H)$ for all graphs $K$, where $\hom(G_{1},G_{2})$ denotes the number of homomorphisms from $G_{1}$ to $G_{2}$. Dvořák showed that it suffices to count homomorphisms from all $2$-degenerate graphs $K$. On the other hand, for several interesting graph classes $\mathcal{M}$, it has been shown that there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. Most such classes are minor-closed classes and Roberson conjectured that every proper minor-closed and union-closed graph class $\mathcal{M}$ has the property of there existing non-isomorphic graphs that are indistinguishable by homomorphism counts from $\mathcal{M}$. There has been an effort to prove Roberson’s conjecture as it is believed that minor-closed classes play a special role in the context of homomorphism indistinguishability.
We show that this special role, if so, must be shared, by proving an analogue of Roberson’s conjecture holds for a rich family of non-minor-closed classes. Namely, we prove that for any proper immersion-closed and union-closed class $\mathcal{M}$, there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. This extends a result of Roberson on homomorphism indistinguishability over bounded degree graphs, and gives an almost full picture since our result cannot be extended in the natural way, that is, by replacing immersions with topological minors, due to a result of Neuen and Seppelt.
Our main result is obtained as a consequence of a colouring result which may be of independent interest: Every graph with a ${t\choose 2}(7t+7)$-oddomorphism admits a $K_{t}$-immersion. We also give an alternative proof of the known result that homomorphism counts from graphs of bounded treewidth does not distinguish non-isomorphic graphs. This again follows from a new result on oddomorphisms: Every graph with a $(t+2)$-oddomorphism has treewidth at least $t+1$. This latter bound is best possible.
@@ 1 Introduction
The most well known of these refinement algorithm is called colour refinement or the one-dimensional Weisfeiler-Leman algorithm. This is a practical graph isomorphism solver and it was shown in [1] that, in a smoothed analysis setting, colour refinement solves the majority of cases of graph isomorphism in near-linear time. While the algorithmic definitions of colour refinement are more useful for actual computation, in this note we will instead consider the homomorphism definition.
Recall that for two graphs $G$, $H$, a graph homomorphism from $G$ to $H$ is a function $f:V(G)\to V(H)$ such that for all $e=xy\in E(G)$, we have $f(x)f(y)\in E(H)$. Let $\hom(G,H)$ denote the number of homomorphisms from $G$ to $H$.
@@
We see from Table 1 that for a given graph class $\mathcal{M}$, it is not obvious what the relation $_{\mathcal{M}}$ is, nor what the complexity of determining if $G=_{\mathcal{M}}H$ is. However, important graph isomorphism tests appear by considering various natural classes $\mathcal{M}$. In Table 1 all the classes $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism are all minor-closed [^1] Given a graph $G$, a graph $H$ is a minor of $G$ if it can be obtained by deleting edges, contracting edges, and deleting isolated vertices. classes, or at least closed under edge contractions. This motivates the following conjecture of Roberson.
@@
If $\mathcal{M}$ is a minor-closed and union-closed class that does not contain all graphs, then there are non-isomorphic graphs $G$ and $H$ such that $G=_{\mathcal{M}}H$.
However, even if this conjecture were true, it would not give the full picture. Indeed if $\mathcal{M}$ is the class of graphs with maximum degree at most $d$, then $_{\mathcal{M}}$ is not the isomorphism relation as shown in [22], and the class of graphs with maximum degree at most $d$ is not minor-closed for any fixed $d\geq 3$. So what is the correct level of generality? A natural starting point to unify bounded degree graphs and minor-closed classes are classes that are closed under topological minors [^2] GG contains $H$ as a topological minor if a subgraph of $G$ contains a subdivision of $H$. Yet recently Neuen and Seppelt [21] showed that homomorphism counting is isomorphism even for the class of graphs with no $K_{5}$ topological minor. The authors of [21] interpret this last result as evidence of the “special role” that minor-closed classes play in demarcating the graph classes $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism. While this special role can only be confirmed by the proof of Conjecture 1.2, we show that the role of immersions is at least as important by proving the following, which is the immersion-analogue of Roberson’s conjecture.
@@
Let $\mathcal{M}$ be a proper immersion-closed and union-closed class. Then there are non-isomorphic graphs $G,H$ such that for all $K\in\mathcal{M}$, we have $\hom(K,G)=\hom(K,H)$.
Immersion is a well-studied [^3] See e.g. [4, 5, 8, 15, 24], as a small sampling. graph order incomparable to the minor order, which nevertheless shares important properties of it: it is a well-quasi-order, as conjectured by Nash-Williams in 1963 [19] and proved by Robertson and Seymour [23]. Moreover, if a graph $G$ contains $H$ as a topological minor, then $G$ contains $H$ both as a minor and as an immersion. In particular, this tells us that the result of Neuen and Seppelt helps to emphasize the importance of Theorem 1.3, since its natural generalisation would be to consider classes closed under topological minors. We now give the definition of immersion.
@@
Let $G$ be a graph, and suppose we have three distinct vertices $u,v,w$ with edges uvuv,vwvw. To split off $uv$ and $vw$ is to delete the edges $uv$ and $vw$, and add the edge $uw$ (possibly adding a parallel edge if necessary). A graph $G$ contains a graph $H$ as an immersion if $H$ can be obtained from a subgraph of $G$ by splitting off pairs of edges and deleting isolated vertices.
Note that for every positive integer $d$, the class of graphs with maximum degree at most $d$ is immersion-closed. Therefore, Theorem 1.3 generalises the result of Roberson [22] that homomorphisms from bounded degree classes does not capture isomorphism. The following further strengthening may be possible and we leave it as an open problem.
@@
To prove Theorem 1.3, we utilize tools developed by Roberson in [22], that essentially reduce the problem to a colouring one. Recall that $N_{G}(v)$ denotes the neighbourhood of $v$ in a graph $G$. The following definitions are introduced in [22].
@@
Of course, it is possible for a vertex to be neither odd nor even.
@@
If $G$ admits a $t$-oddomorphism, we write $G\xrightarrow{\text{odd}}K_{t}$. In Section 3 we prove the following.
@@
While it does not appear at all related, it turns out that Theorem 1.8 and a result in [22] implies Theorem 1.3. We show that this is the case in Section 5. The bound we obtain is not necessarily optimal, and we make the following conjecture.
@@
It might even be possible that $c=1$ can be taken. In fact, Theorem 1.10 below gives that $c=1$ can indeed be taken when $t=2,3$. This theorem, moreover, is a tight lower bound on the treewidth of graphs that have a $t$-oddomorphism. (We recall the definition of treewidth in Section 2.3.)
@@
If $f:G\xrightarrow{\text{odd}}K_{t+2}$, then $\operatorname{tw}(G)\geq t+1$.
@@
The $t+1$ bound given here is best possible since $K_{t+2}\xrightarrow{\text{odd}}K_{t+2}$ and $\operatorname{tw}(K_{t+2})=t+1$. Theorem 1.10 also gives an alternative proof that homomorphism counts from graphs of bounded treewidth graph do not fully capture isomorphism. However, as we have mentioned above, homomorphism counts from bounded treewidth graphs is exactly described by the $k$-dimensional Weisfeiler-Leman algorithm, which is already known to not fully capture isomorphism [20].
The rest of the paper is structured as follows. In Section 2.3 we give some definitions, state a known key result, and prove some preliminary lemmas about oddomorphisms. In Section 3, we prove Theorem 1.8. In Section 4, we prove Theorem 1.10. Finally, in Section 5, we prove Theorem 1.3. We devote the rest of this section to sketch the proofs of our oddomorphism results.
@@
To prove Theorem 1.8 the idea is as follows. We observe first that if $f$ is a ${t\choose 2}(7t+7)$-oddomorphism, then if there are any $2$-coloured cycles, deleting the edges of the cycle preserves having an oddomorphism. We therefore may assume that the union of any two colour classes is a forest. Next we observe if the union of two colour classes has a path from an $f$-odd vertex in one colour to an $f$-odd vertex in another colour, then we can split off the entire path and still have an ${t\choose 2}(7t+7)$-oddomorphism. Thus the structure of a minimal counterexample is that between any two colour classes, say $C_{i}$ and $C_{j}$, in a component that contains an $f$-odd vertex $v$ say in $C_{i}$, we have that either $v$ is adjacent to an $f$-odd vertex in $C_{j}$, or there is a $2$-coloured path $P$ in $G[C_{i},C_{j}]$ with endpoints $v$ and another $f$-odd vertex in $C_{i}$. In a systematic way, we split off all these paths, and get a new graph $G^{\prime}$ (which need not have an ${t\choose 2}(7t+7)$-oddomorphism).
Utilizing the $f$-odd properties of oddomorphisms, there are two possibilities for $G^{\prime}$. Either there will be two vertices given the same colour $x,y$ which have at least $t\choose 2$} parallel edges between them, or the graph obtained after removing parallel edges has minimum degree at least $7t+7$. In the latter case, a theorem of Gauthier, Le and Wollan [11] gives the $K_{t}$-immersion.
@@
In the former case, we have two vertices with the same colour that have at least $t\choose 2$} edge disjoint $2$-coloured paths between them, and identifying these vertices, deleting the $2$-coloured paths, and removing the symmetric difference of the neighbourhoods, gives a smaller graph with an ${t\choose 2}(7t+7)$-oddomorphism, and we show that the $K_{t}$-immersion of the smaller graph lifts to an immersion of the original graph.
@@
To prove Theorem 1.10 the idea is the following. Suppose we have a $(t+2)$-oddomorphism and our graph has treewidth at most $t$, and that we have a vertex minimal counterexample. Take an optimal tree decomposition which we can assume has some nice “smooth” properties, and root the tree decomposition at some vertex. By using a breadth first search from the root, find an odd vertex which appears last (so we count from the root each the first time each odd vertex appears in a bag, and we take the last odd vertex to appear). Since we have a $(t+2)$-oddomorphism, and each bag has size at most $t+1$, this odd vertex will have a neighbour in a bag farther away from the root, and further the colour of this neighbour will not appear in its current bag. Then by looking at a Kempe chain between the two colours of the odd vertex and this other vertex lower in the tree, we will find two vertices with the same colour that we can identify and keep the treewidth the same, and preserve the oddomorphism. In this way we find a smaller counterexample, which is a contradiction.
@@ 2.2 Definition of treewidth
Let $G$ be a graph. A tree decomposition of $G$ is a pair $\left(T,\{\beta_{t}\}_{t\in V(T)}\right)$ where $T$ is a tree and each $\beta_{t}\subseteq V$ (called a bag) such that:
@@
    for every edge $uv\in E(G)$, there exists $t\in V(T)$ with $uv\in E(G[\beta_{t}])$;
@@
    for every vertex $v\in V(G)$, the set $\{t\in V(T)\mid v\in\beta_{t}\}$ induces a connected subtree of $T$.
The width of a tree decomposition $\left(T,\{\beta_{t}\}_{t\in V(T)}\right)$ is $\max_{t\in V(T)}|\beta_{t}|-1$. The treewidth of a graph $G$, denoted $\operatorname{tw}(G)$, is the minimum width over all tree decompositions of $G$.
@@ 2.3 Preliminaries about oddomorphisms
3 Large oddomorphism implies large complete immersion
For ease, given an $(x,y)$-path $P$, we say that to split a path $P$ is to iteratively split off edges of $P$ to create the edge $xy$.
@@ 3 Large oddomorphism implies large complete immersion
By the minimality of $G$, $G^{*}$ has a $K_{t}$-immersion. We claim that this implies that $G$ has a $K_{t}$-immersion. To see this, we use the definition of immersions given in Section 2.3. Let $t_{1},\ldots,t_{t}$ be the branch vertices of the $K_{t}$-immersion in $G^{*}$ and, for each $1\leq i<j\leq t$, let $Q_{ij}$ be the path from $t_{i}$ to $t_{j}$ (such that $Q_{ij}$’s are mutually edge-disjoint). Observe that if $Q_{ij}$ uses the merged vertex more than once, $Q_{ij}$ in fact contains a cycle. Thus each path $Q_{ij}$ uses the merged vertex at most once. Therefore we can extend this immersion to an immersion of $G$ in the following manner. For each $Q_{ij}$ that uses the merged vertex, we use one of the paths $P_{1},\ldots,P_{{t\choose 2}}$ to extend $Q_{i,j}$ to a path in $G$. As there are at most $t\choose 2$} paths $Q_{ij}$, we can do this such that each $P_{i}$ is used at most once, and thus we have a $K_{t}$-immersion in $G$.

### v2

# Homomorphism counting from a proper immersion-closed class is not isomorphism

@@ Abstract
Lovász proved that two graphs $G$ and $H$ are isomorphic if $\hom(K,G)=\hom(K,H)$ for all graphs $K$, where $\hom(G_{1},G_{2})$ denotes the number of homomorphisms from $G_{1}$ to $G_{2}$. Dvořák showed that it suffices to count homomorphisms from all $2$-degenerate graphs $K$. On the other hand, for several interesting graph classes $\mathcal{M}$, it has been shown that there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. Most such classes are minor-closed and Roberson conjectured that every proper minor-closed graph class $\mathcal{M}$ has the property that there exist non-isomorphic graphs that are indistinguishable by homomorphism counts from $\mathcal{M}$. There has been an effort to prove Roberson’s conjecture as it is believed that minor-closed classes play a special role in demarcating the graph classes that satisfy this property.
We show that this special role, if it exists, must be shared, by proving an analogue of Roberson’s conjecture for a rich family of non-minor-closed classes. Namely, we prove that for any proper immersion-closed graph class $\mathcal{M}$, there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. This extends a result of Roberson on homomorphism indistinguishability over bounded degree graphs, and cannot be extended in the natural way by replacing immersions with topological minors due to a result of Neuen and Seppelt.
Our main result is obtained as a consequence of a colouring result which may be of independent interest: Every graph with a ${t\choose 2}(7t+7)$-oddomorphism admits a $K_{t}$-immersion. We also give a shorter proof of a result of Neuen that every graph with a $t$-oddomorphism has treewidth at least t1t-1, which implies that homomorphism counts from graphs of bounded treewidth does not distinguish non-isomorphic graphs.
@@ 1 Introduction
The most well known of these refinement algorithm is called colour refinement or the one-dimensional Weisfeiler-Leman algorithm. This refinement algorithm is used in practical graph isomorphism solvers and it was shown in [1] that, in a smoothed analysis setting, colour refinement solves the majority of cases of graph isomorphism in near-linear time. While the algorithmic definitions of colour refinement are more useful for actual computation, in this note we will instead consider the homomorphism definition.
Recall that for two graphs $G$ and $H$, a graph homomorphism from $G$ to $H$ is a function $f:V(G)\to V(H)$ such that for all $e=xy\in E(G)$, we have $f(x)f(y)\in E(H)$. Let $\hom(G,H)$ denote the number of homomorphisms from $G$ to $H$.
@@
We see from Table 1 that for a given graph class $\mathcal{M}$, it is not obvious what the relation $_{\mathcal{M}}$ is, nor what the complexity of determining if $G=_{\mathcal{M}}H$ is. However, important graph isomorphism tests appear by considering various natural classes $\mathcal{M}$. In Table 1, every class $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism is contained in a minor-closed [^1] Given a graph $G$, a graph $H$ is a minor of $G$ if it can be obtained by deleting edges, contracting edges, and deleting isolated vertices. A graph class $\mathcal{M}$ is minor-closed if every minor of every graph in $\mathcal{M}$ is also in $\mathcal{M}$. class. This motivates the following conjecture of Roberson. A graph class $\mathcal{M}$ is proper if it does not contain all graphs and it is union-closed if the disjoint union of any two graphs in $\mathcal{M}$ is also in $\mathcal{M}$.
@@
If $\mathcal{M}$ is a proper graph class that is minor-closed and union-closed, then there are non-isomorphic graphs $G$ and $H$ such that $G=_{\mathcal{M}}H$.
However, even if this conjecture were true, it would not give the full picture. Indeed if $\mathcal{M}$ is the class of graphs with maximum degree at most $d$, then $_{\mathcal{M}}$ is not the isomorphism relation as shown in [23], and the class of graphs with maximum degree at most $d$ is not minor-closed for any fixed $d\geq 3$. So what is the correct level of generality? A natural starting point to unify bounded degree graphs and minor-closed classes are classes that are closed under topological minors [^2] GG contains $H$ as a topological minor if a subgraph of $G$ contains a subdivision of $H$. Yet recently Neuen and Seppelt [22] showed that homomorphism counting is isomorphism even for the class of graphs with no $K_{5}$ topological minor. The authors of [22] interpret this last result as evidence of the “special role” that minor-closed classes play in demarcating the graph classes $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism. However, we show that the role of immersions is at least as important as minors by proving the following immersion-analogue of Roberson’s conjecture.
@@
If $\mathcal{M}$ is a proper graph class that is immersion-closed and union-closed, then there are non-isomorphic graphs $G$ and $H$ such that for all $K\in\mathcal{M}$, we have $\hom(K,G)=\hom(K,H)$.
While a proper immersion-closed graph class $\mathcal{M}$ may not be union-closed, it is contained in a proper graph class $\mathcal{M}^{\prime}$ that is immersion-closed and union closed. In particular, if $G$ and $H$ are non-isomorphic graphs such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}^{\prime}$, then we also have $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. Hence, Theorem 1.3 holds for every proper immersion-closed graph class.
Immersion is a well-studied [^3] See e.g. [4, 6, 9, 16, 25], as a small sampling. graph order incomparable to the minor order, which nevertheless shares important properties of it: it is a well-quasi-order, as conjectured by Nash-Williams in 1963 [20] and proved by Robertson and Seymour [24]. Moreover, if a graph $G$ contains $H$ as a topological minor, then $G$ contains $H$ both as a minor and as an immersion. In particular, this tells us that the result of Neuen and Seppelt helps to emphasize the importance of not only Conjecture 1.2 but Theorem 1.3 as well, since graph classes closed under topological minors are natural generalizations of both immersion-closed classes and minor-closed classes. We now give the definition of immersion.
@@
Let $G$ be a graph, and suppose we have three distinct vertices $u,v,w$ with edges uvuv,vwvw. To split off $uv$ and $vw$ is to delete the edges $uv$ and $vw$, and add the edge $uw$ (possibly adding a parallel edge if necessary). A graph $G$ contains a graph $H$ as an immersion if $H$ can be obtained from a subgraph of $G$ by splitting off pairs of edges and deleting isolated vertices. A graph class $\mathcal{M}$ is immersion-closed if for every $G\in\mathcal{M}$ and every graph $H$, if $G$ contains $H$ as an immersion, then $H\in\mathcal{M}$.
Note that for every positive integer $d$, the class of graphs with maximum degree at most $d$ is immersion-closed. Therefore, Theorem 1.3 generalizes the result of Roberson [23] that homomorphisms from bounded degree classes does not capture isomorphism. The following further strengthening may be possible and we leave it as an open problem.
@@
To prove Theorem 1.3, we utilize tools developed by Roberson in [23], which essentially reduce the problem to a colouring one. Recall that $N_{G}(v)$ denotes the neighbourhood of $v$ in a graph $G$. The following definitions are introduced in [23].
@@
Of course, it is possible for a vertex to be neither $f$-odd nor $f$-even.
@@
If $G$ admits a $t$-oddomorphism $f$, we say that $f$ is an oddomorphism from $G$ to $K_{t}$ and write $G\xrightarrow{\text{odd}}K_{t}$. In Section 3 we prove the following.
@@
While it does not appear at all related, it turns out that Theorem 1.8 and a result in [23] implies Theorem 1.3. We show that this is the case in Section 5. The bound in Theorem 1.8 is not necessarily optimal, and we make the following conjecture.
@@
It might even be possible that $c=1$ can be taken. When $t\leq 4$, we can indeed take $c=1$ due to a tight lower bound on the treewidth $\operatorname{tw}(G)$ of graphs $G$ admitting a $t$-oddomorphism below. (We recall the definition of treewidth in Section 2.) While Theorem 1.10 follows from the proof of [21, Theorem 1.2] by Neuen, [^4] Neuen proves a more general result that if $G$ admits an oddomorphism to any graph $H$ (see Definition 5.4), then $\operatorname{tw}(G)\geq\operatorname{tw}(H)$. we give a short elementary proof of this fact.
@@
If $f:G\xrightarrow{\text{odd}}K_{t}$, then $\operatorname{tw}(G)\geq t-1$.
@@
The t1t-1 bound given here is best possible since $K_{t}\xrightarrow{\text{odd}}K_{t}$ and $\operatorname{tw}(K_{t})=t-1$. Theorem 1.10 also gives an alternative proof that homomorphism counts from graphs of bounded treewidth graph do not fully capture isomorphism. However, as we have mentioned above, homomorphism counts from bounded treewidth graphs is exactly described by the $k$-dimensional Weisfeiler-Leman algorithm, which is already known to not fully capture isomorphism ([5]). The rest of the paper is structured as follows. In Section 2 we give some definitions, state a known key result, and prove some preliminary lemmas about oddomorphisms. In Section 3, we prove Theorem 1.8. In Section 4, we prove Theorem 1.10. Finally, in Section 5, we prove Theorem 1.3. We devote the rest of this section to sketch the proofs of our oddomorphism results.
@@
To prove Theorem 1.8, we observe first that oddomorphisms are preserved when deleting a $2$-coloured cycle and when splitting off a $2$-coloured path between two $f$-odd vertices of different colours. Thus, in a minimal counterexample, any two colour classes induce a forest in which each component has $f$-odd vertices of exactly one colour. We further split off 2-coloured paths between $f$-odd vertices in these components to obtain a new graph $G^{\prime}$. Note that the oddomorphism is no longer preserved. Nonetheless, if $f$ is a ${t\choose 2}(7t+7)$-oddomorphism of $G$, then either $G^{\prime}$ contains two vertices of same colour with at least $t\choose 2$} parallel edges between them, or the graph obtained from $G^{\prime}$ by removing parallel edges has minimum degree at least $7t+7$.
@@
In the former case, we identify the two vertices in $G$, delete the $2$-coloured paths between them, and take the symmetric difference of their neighbourhoods. This gives a smaller graph with an ${t\choose 2}(7t+7)$-oddomorphism, and we show that the $K_{t}$-immersion of the smaller graph lifts to an immersion of the original graph. In the latter case, a theorem of Gauthier, Le and Wollan [12] gives the $K_{t}$-immersion.
@@
To prove Theorem 1.10, suppose $G$ has a $t$-oddomorphism $f$ and has treewidth at most t2t-2, and that $G$ is vertex minimal with respect to these properties. Take a tree decomposition of minimum width and root the tree decomposition at some vertex. We choose a non-leaf bag $\beta_{v}$ containing an $f$-odd vertex $w$ such that no other $f$-odd vertex appears farther from the root than $w$. Since $f$ is a $t$-oddomorphism and each bag has size at most t1t-1, $w$ has a neighbour $w^{\prime}$ in a bag farther away from the root whose colour does not appear in $\beta_{v}$. By looking at the Kempe chain between these two colours, we find two vertices with the same colour in $\beta_{v}$ that we can identify without increasing the treewidth while preserving the oddomorphism, contradicting the minimality of $G$.
@@ 2.2 Definition of treewidth
Let $G$ be a graph. A tree decomposition of $G$ is a pair $\left(T,\{\beta_{x}\}_{x\in V(T)}\right)$ where $T$ is a tree and each $\beta_{x}\subseteq V$ (called a bag) such that:
@@
    for every edge $uv\in E(G)$, there exists $x\in V(T)$ with $uv\in E(G[\beta_{x}])$;
@@
    for every vertex $v\in V(G)$, the set $\{x\in V(T)\mid v\in\beta_{x}\}$ induces a connected subtree of $T$.
The width of a tree decomposition $\left(T,\{\beta_{x}\}_{x\in V(T)}\right)$ is $\max_{x\in V(T)}|\beta_{x}|-1$. The treewidth of a graph $G$, denoted $\operatorname{tw}(G)$, is the minimum width over all tree decompositions of $G$.
@@ 2.3 Preliminaries about oddomorphisms
3 Large oddomorphism implies large clique immersion
In this section we prove Theorem 1.8. Given an $(x,y)$-path $P$, we say that to split a path $P$ is to iteratively split off edges of $P$ to create the edge $xy$.
@@ 3 Large oddomorphism implies large clique immersion
By the minimality of $G$, $G^{*}$ has a $K_{t}$-immersion. We claim that this implies that $G$ has a $K_{t}$-immersion. To see this, we use the definition of immersions given in Section 2. Let $t_{1},\ldots,t_{t}$ be the branch vertices of the $K_{t}$-immersion in $G^{*}$ and, for each $1\leq i<j\leq t$, let $Q_{ij}$ be the path from $t_{i}$ to $t_{j}$ in $G^{*}$ (such that $Q_{ij}$’s are mutually edge-disjoint). Observe that if $Q_{ij}$ uses the merged vertex more than once, $Q_{ij}$ in fact contains a cycle. Thus each path $Q_{ij}$ uses the merged vertex at most once. Therefore we can extend this immersion to an immersion of $G$ in the following manner. For each $Q_{ij}$ that uses the merged vertex, we use one of the paths $P_{1},\ldots,P_{{t\choose 2}}$ to extend $Q_{i,j}$ to a path in $G$. As there are at most $t\choose 2$} paths $Q_{ij}$, we can do this such that each $P_{i}$ is used at most once, and thus we have a $K_{t}$-immersion in $G$.